# Signed Higher-Order Interactions for Brain Disorder Diagnosis via Multi-Channel Transformers

### v2

@@ 4.4.3 The important brain regions and interactions of brain regions
Alzheimer’s disease (AD). From the Fig. 12, we observed a significant overlap between the brain regions of interest and those implicated in critical interactions, particularly involving the left caudate nucleus. This may be attributed to the damage sustained by relevant brain regions during the progression of Alzheimer’s disease, which subsequently leads to weakened connectivity between brain regions. The ten ROIs with the highest attention scores include: Right Caudate (CAU.R), Left Hippocampus (HIP.L), Right Parahippocampal Gyrus (PHG.R), Left Amygdala (AMYG.L), Left Olfactory Cortex (OLF.L), Left Middle Temporal Gyrus (TPOmid.L), Right Olfactory Cortex (OLF.R), Left Parahippocampal Gyrus (PHG.L), Left Caudate (CAU.L), and Right Heschl’s Gyrus (HES.R). These regions are primarily associated with functions such as memory (Hippocampus, Parahippocampal gyrus), emotion (Amygdala), olfaction (Olfactory cortex), motor control (Caudate nucleus), and auditory processing (Heschl’s gyrus). Specifically, the Hippocampus and Parahippocampal gyrus have been widely confirmed to be associated with Alzheimer’s disease (AD) (Bv and Agrawal, 2022; Echávarri et al., 2010), as they are involved in brain memory functions. The Amygdala has also been widely demonstrated to undergo atrophy during the progression of diseases (Zidan et al., 2019). And the Caudate, due to its crucial role in regulating motor control, has significant associations with AD (Zhi et al., 2024). Ye et al. also confirmed the association between the Middle Temporal Gyrus and AD (Ye et al., 2019). The Olfactory Cortex and the Heschl’s Gyrus currently lack direct evidence linking them to Alzheimer’s disease. However, they may be associated with the loss of hearing and smell in Alzheimer’s patients, which suggests that this finding obtained by HOI-brain may provide new insights into the pathology of Alzheimer’s disease. In addition, we observed that most of the top ten ROI interactions with the highest attention scores revolve around the Right Caudate, suggesting that the Right Caudate may serve as a hub node in the brain network and is disrupted during the progression of Alzheimer’s disease, which is consistent with the literature (Zhi et al., 2024).
@@
Parkinson’s syndrome (PD). As illustrated in Fig. 13, the ten regions of interest (ROIs) with the highest attention scores are: Right Precentral Gyrus (PCG.R), Left Medial Orbitofrontal Cortex (ORBsupmed.L), Right Inferior Parietal Lobule (IPL.R), Left Amygdala (AMYG.L), Right Triangular Part of the Inferior Frontal Gyrus (IFGtriang.R), Right Superior Parietal Gyrus (SPG.R), Right Inferior Orbitofrontal Cortex (ORBinf.R), Right Thalamus (THA.R), Right Superior Temporal Gyrus (STG.R), and Right Opercular Part of the Inferior Frontal Gyrus (IFGoperc.R). These regions are predominantly implicated in motor execution, reward processing, executive control, affective regulation, and sensorimotor integration—domains known to be disrupted in Parkinson’s disease (PD). Specifically, the Precentral Gyrus (PCG.R), a key component of the primary motor cortex, exhibits marked hypometabolism and altered functional connectivity in PD (Wang et al., 2024b). The left medial orbitofrontal cortex (ORBsupmed.L) is related to decision-making performance in Parkinson’s disease (Kobayakawa et al., 2016). Moreover, compared with healthy controls (HCs), patients with Parkinson’s disease (PD) showed a globally reduced structural-functional connectivity decoupling in the Inferior and Medial Orbitofrontal Cortices (ORBinf.R and ORBsupmed.L) (Zou et al., 2024). The Amygdala (AMYG.L), central to emotional processing and reward evaluation, demonstrates reduced volume and disrupted connectivity within limbic-striatal circuits in PD, which may underlie affective dysregulation and apathy (Yoshimura et al., 2005). The Thalamus (THA.R), a vital relay hub within cortico-basal ganglia-thalamo-cortical loops, is frequently implicated in PD pathophysiology due to its role in motor and non-motor symptomatology, including tremor and cognitive fluctuations (Zirh et al., 1998). The Triangular and Opercular parts of the Inferior Frontal Gyrus (IFGtriang.R and IFGoperc.R), critical for executive control and language processing, demonstrate impaired activation and connectivity in PD, aligning with cognitive inflexibility and verbal fluency deficits (Li et al., 2020). It is worth noting that although there is currently no direct literature indicating a pathological association between the superior temporal gyrus (STG.R) and Parkinson’s disease, its involvement in auditory processing and social cognition may indirectly contribute to communication difficulties and social withdrawal symptoms in patients with Parkinson’s disease. Crucially, unlike ADNI and ASD, the ”core hubs” in the interaction network—PCG.R, STG.R, IFGtriang.R, and ANG.R—did not rank among the top positions in the ”key brain importance” list. Conversely, the seven regions listed as ”key brain importance,” such as ACG.R, TPOmid.R, and CAU.L, were almost entirely absent from the top 20 strongest interactions. Together, these findings highlight two distinct perspectives on ”importance” in brain: ”global centrality” versus ”local strong coupling.”
@@
Autism Spectrum Disorder (ASD). As illustrated in Fig. 14, the ten regions of interest (ROIs) with the highest attention scores are: Right Inferior Temporal Gyrus (ITG.R), Left Supramarginal Gyrus (SMG.L), Left Inferior Temporal Gyrus (ITG.L), Left Triangular Part of the Inferior Frontal Gyrus (IFGtriang.L), Left Caudate (CAU.L), Right Caudate (CAU.R), Left Middle Temporal Gyrus (TPOmid.L), Right Middle Temporal Gyrus (TPOmid.R), Right Anterior Cingulate Gyrus (ACG.R), and Right Rolandic Operculum (ROL.R). These regions are predominantly implicated in social cognition, language processing, executive control, and sensorimotor integration. Specifically, the Inferior Temporal Gyrus (ITG) has been repeatedly reported to exhibit atypical activation and connectivity patterns in individuals with ASD, reflecting its crucial role in face perception and semantic processing (Kim et al., 2021). The Supramarginal Gyrus, a core component of the mirror-neuron system, has also been shown to demonstrate aberrant functional connectivity in ASD, which may contribute to deficits in empathy and social interaction (Bi et al., 2025). The Caudate nucleus, given its involvement in reward processing and cognitive flexibility, has been associated with restricted and repetitive behaviors characteristic of ASD (Adorjan et al., 2017). Furthermore, the Middle Temporal Gyrus and the Anterior Cingulate Gyrus have been implicated in the pathophysiology of ASD through their roles in theory-of-mind and conflict monitoring, respectively (Zhu et al., 2022; Xu et al., 2019). The Rolandic Operculum, although less frequently discussed, is closely linked to auditory and sensorimotor integration, the disruption of which may underlie sensory hypersensitivity commonly observed in ASD (Zhang et al., 2025). Notably, the majority of the top ten ROI interactions with the highest attention scores center on ITG.R, suggesting that this region may also function as a pivotal hub within the brain network and is selectively disrupted during the progression of ASD. This finding aligns with recent literature highlighting the central role of the ITG in the neurodevelopmental trajectory of autism (Kim et al., 2021).
@@ 4.4.4 The higher-order organization that differ between patients and healthy individuals
Next, we further investigate whether the higher-order interactions among key brain regions identified by HOI-Brain show significant differences during disease progression, which could offer additional therapeutic insights for brain disease research. For ease of calculation, we selected four out of the top ten key brain regions identified for each disease. Using our proposed Multiplication of Temporal Derivatives (MTD)-based co-fluctuation calculation method, we computed the quadruplet interactions among the relevant brain regions and subsequently compared the group differences. As shown in Fig. 15, 16, 17, compared to other brain regions, although these brain regions all demonstrate significant importance in the diagnosis of Alzheimer’s disease (AD) or Autism Spectrum Disorder (ASD)—indicating that they may have sustained varying degrees of damage—their higher-order interactions exhibit notable differences across various stages of disease progression. This suggests a dissociation between higher-order interactions in the brain and the functions of the individual brain regions themselves, which is consistent with the literature (Wang et al., 2019). In addition, we observed that during the transition from healthy individuals to those with Alzheimer’s disease (AD) or autism spectrum disorder (ASD), the positive higher-order interactions between relevant brain regions gradually weaken, indicating a progressive impairment of higher-order synergistic functions among brain regions. On the other hand, the negative higher-order interactions between these brain regions gradually strengthen, reflecting an increasing degree of internal functional disruption in the brain. This phenomenon is consistent with findings reported in previous literature (Santoro et al., 2024). However, this phenomenon is exactly the opposite in Parkinson’s disease, which may be due to the fact that Parkinson’s patients, because of their motor impairments, are prone to hallucinations. Hallucinations in Parkinson’s disease may result from excessive influence of higher-order brain regions on early sensory processing, which is consistent with the enhanced functional integration between sensory and higher-order networks (Tan et al., 2023).
@@
Alzheimer’s disease (AD). As shown in Fig. 15, in particular, the quadruplet interactions among CAU.R, HIP.L, PHG.R, and AMYG.L; the quadruplet interactions among OLF.L, HIP.L, PHG.R, and AMYG.L; and the quadruplet interactions among CAU.R, OLF.L, PHG.R, and AMYG.L, which involve the cortico-striatal-olfactory circuit, exhibit significant inter-group differences during the progression from cognitively normal (CN) to mild cognitive impairment (MCI), with p-values <0.05<0.05< 0.05. However, these differences become less pronounced during the progression from MCI to AD, with p-values 0.05$\geq$ 0.05≥ 0.05. This suggests that these higher-order interactions may serve as potential biomarkers for the early diagnosis of Alzheimer’s disease. In contrast, the quadruplet interactions among CAU.R, TPOmid.L, HIP.L, and AMYG.L, which involves the ’social-reward-emotion-memory’ integrated circuit, shows significant inter-group differences throughout the entire AD progression, with p-values <0.001<0.001< 0.001. This indicates that these higher-order interactions may be persistently disrupted across the entire disease course, playing a critical role in AD pathogenesis.
@@
Parkinson’s disease (PD). As shown in Fig. 16, the quadruplet interactions among PCG.R, ORBsupmed.L, IPL.R, and SPG.R; the interactions among PCG.R, ORBsupmed.L, IPL.R, and IFGoperc.R; and those among PCG.R, ORBsupmed.L, IPL.R, and STG.R, which collectively engage the dorsal attention network (DAN) and fronto-parietal executive control circuits, exhibit significant inter-group differences ($p<0.05$) between healthy controls and early-stage PD patients. These findings suggest that these functional connectivity patterns may serve as potential early biomarkers for cognitive-motor integration deficits characteristic of the prodromal stages of PD. In addition, the quadruplet interactions involving PCG.R, ORBsupmed.L, THA.R, and IFGoperc.R, which map onto the mesocortical limbic-striatal-thalamic loop, show inter-group differences ($p<0.001$). This persistent disruption aligns with the progressive degeneration of dopaminergic pathways and thalamocortical relay integrity, implicating these interactions in both motor dysfunction (e.g., bradykinesia) and nonmotor symptoms (e.g., apathy, executive dysfunction) throughout the pathogenesis of PD-related motor and cognitive impairment.
@@
Autism Spectrum Disorder (ASD). From Fig. 17, we observed that the quadruplet interactions among the top five key brain regions—including the interactions between ITG.R, SMG.L, ITG.L, and IFGtriang.L; between SMG.L, ITG.L, IFGtriang.L, and CAU.L; between ITG.R, ITG.L, IFGtriang.L, and CAU.L; between ITG.R, SMG.L, IFGtriang.L, and CAU.L; and between ITG.R, SMG.L, ITG.L, and CAU.L, which involves a ”visual-auditory-semantic-action-control” closed-loop system. These interactions exhibited significant inter-group differences throughout the entire progression of Autism Spectrum Disorder (ASD). Such quadruplet interactions may serve as potential biomarkers for investigating tool-use deficits, semantic control impairments, and action-language disconnection in social communication associated with autism. However, the quadruplet interactions among ITG.L, IFGtriang.L, CAU.L, and CAU.R show no significant inter-group differences throughout the entire AD progression, with p-values 0.05$\geq$ 0.05≥ 0.05. This suggests a dissociation between higher-order brain interactions and the functions of individual brain regions themselves.
@@ 5.1 The model
In this study, we focused on the importance of signed higher-order (group) interactions in fMRI signals, which were inferred using a recently developed topological approach (Hyekyoung et al., 2011). Our proposed HOI-Brain incorporates: (i) a novel higher-order interaction representation strategy via Multiplication of Temporal Derivatives (MTD) to quantify dynamic functional co-fluctuations of group ROIs; (ii) a novel higher-order feature extraction method that links higher-order interactions (HOIs) in the brain to higher-order organizational structures, focusing on quadruplet-level interaction signatures and two-dimensional void descriptors extracted through Persistent Homology; (iii) a pioneering effort to distinguish between positively and negatively synergistic higher-order interactions in the brain; (iv) a novel multi-channel brain network Transformer that integrates lower-order edge features with higher-order topological invariants. This new framework emphasizes the potential of quadruplet or higher-order region interactions — representing critical advancements for precision medicine, a deeper understanding of neurological disorders, and broader contributions to neuroimaging research. It demonstrates superior prediction accuracy for classifying Alzheimer’s Disease (AD), Parkinson’s Disease (PD), and Autism Spectrum Disorder (ASD) compared to traditional machine learning models, graph neural network (GNN)-based models, Transformer-based models, hypergraph neural network (HGNN)-based models, and the Persistent Homology (PH)-based models. Additionally, ablation study also indicates that the effectiveness of quadruplet interactions compared to triplet interactions, the effectiveness of distinguishing between positively and negatively synergistic higher-order interactions, and the effectiveness of different components in model.
@@
Furthermore, an investigation was conducted into the most discriminative neural patterns using an attention mechanism. This investigation revealed that quadruplet-level interaction signatures exhibit substantially greater importance than two-dimensional void descriptors and edge features in both the ADNI and ABIDE datasets. It is noteworthy that negative synergistic quadruplets tend to exhibit a higher degree of significance in comparison to their positive synergistic counterparts. Furthermore, positive synergistic voids have been observed to exceed negative synergistic voids in terms of their impact. This novel observation may offer critical insights into the pathological mechanisms underlying brain disorders. In order to provide further validation of these findings, inter-group comparisons were performed of signed higher-order topological signatures in the ADNI dataset. The findings demonstrated a substantial discrepancy in the quantity of quadruplet-level interaction signatures among the groups ($p<0.001$) in comparison to void descriptors. Furthermore, positive voids demonstrated more pronounced inter-group disparities than negative voids. The findings suggest that disease progression may primarily disrupt higher-order neural interactions, which subsequently influence high-dimensional hole structures. An intriguing pattern emerged in the comparative analysis of cognitively normal (CN) individuals and those with Alzheimer’s disease (AD). CN individuals appear to rely more on positive information for whole-brain topological organisation, whereas AD patients show a greater dependence on negative information. Furthermore, longitudinal trends indicate that the number of positive quadruplets gradually declines with disease progression, while negative quadruplets follow a biphasic trajectory – decreasing initially and then increasing. This suggests that during the transition from CN to mild cognitive impairment (MCI), there is a decline in both positive and negative synergistic interactions. However, as the condition progresses from MCI to AD, there is an increasing adoption of negative synergy by neural networks, alongside a suppression of excitability. This may be a compensatory mechanism employed by the brain to preserve functional integrity across neural modules.
@@
In the context of investigating the significant brain regions and their interactions in the context of brain diseases, the utilisation of HOI-Brain facilitates the identification of these regions from a more comprehensive perspective, a departure from the low-level perspective that has been adopted in previous studies (Yan et al., 2019; Li et al., 2021). The findings, which function as potential biomarkers, are consistent with a substantial corpus of extant literature (Zidan et al., 2019; Zhi et al., 2024; Ye et al., 2019; Kim et al., 2021). Furthermore, they furnish novel insights with regard to the diagnosis of brain diseases. The right caudate has been identified as a potential hub node within the brain network. Previous literature has indicated that disruptions in this region are consistent with the progression of Alzheimer’s disease (Zhi et al., 2024). In a similar manner, the right inferior temporal gyrus (ITG.R) has been observed to function as an additional pivotal hub within the brain network. It has been demonstrated that this region is selectively disrupted during the progression of Autism Spectrum Disorder (ASD) (Kim et al., 2021).

### v3

@@ 4.4.3 The important brain regions and interactions of brain regions
Alzheimer’s disease (AD). From the Fig. 12, we observed a significant overlap between the ROIs and those implicated in critical interactions, particularly involving the left caudate nucleus. This may be attributed to the damage sustained by relevant brain regions during the progression of Alzheimer’s disease, which subsequently leads to weakened connectivity between brain regions. The ten ROIs with the highest attention scores include: Right Caudate (CAU.R), Left Hippocampus (HIP.L), Right Parahippocampal Gyrus (PHG.R), Left Amygdala (AMYG.L), Left Olfactory Cortex (OLF.L), Left Middle Temporal Gyrus (TPOmid.L), Right Olfactory Cortex (OLF.R), Left Parahippocampal Gyrus (PHG.L), Left Caudate (CAU.L), and Right Heschl’s Gyrus (HES.R). These regions are primarily associated with functions such as memory (Hippocampus, Parahippocampal gyrus), emotion (Amygdala), olfaction (Olfactory cortex), motor control (Caudate nucleus), and auditory processing (Heschl’s gyrus). Specifically, the Hippocampus and Parahippocampal gyrus have been widely confirmed to be associated with AD (Bv and Agrawal, 2022; Echávarri et al., 2010), as they are involved in brain memory functions. The Amygdala has also been widely demonstrated to undergo atrophy during the progression of diseases (Zidan et al., 2019). And the Caudate, due to its crucial role in regulating motor control, has significant associations with AD (Zhi et al., 2024). Ye et al. also confirmed the association between the Middle Temporal Gyrus and AD (Ye et al., 2019). The Olfactory Cortex and the Heschl’s Gyrus currently lack direct evidence linking them to Alzheimer’s disease. However, they may be associated with the loss of hearing and smell in Alzheimer’s patients, which suggests that this finding obtained by HOI-brain may provide new insights into the pathology of Alzheimer’s disease. In addition, we observed that most of the top ten ROI interactions with the highest attention scores revolve around the Right Caudate, suggesting that the Right Caudate may serve as a hub node in the brain network and is disrupted during the progression of Alzheimer’s disease, which is consistent with the literature (Zhi et al., 2024).
@@
Parkinson’s syndrome (PD). As illustrated in Fig. 13, the ten ROIs with the highest attention scores are: Right Precentral Gyrus (PCG.R), Left Medial Orbitofrontal Cortex (ORBsupmed.L), Right Inferior Parietal Lobule (IPL.R), Left Amygdala (AMYG.L), Right Triangular Part of the Inferior Frontal Gyrus (IFGtriang.R), Right Superior Parietal Gyrus (SPG.R), Right Inferior Orbitofrontal Cortex (ORBinf.R), Right Thalamus (THA.R), Right Superior Temporal Gyrus (STG.R), and Right Opercular Part of the Inferior Frontal Gyrus (IFGoperc.R). These regions are predominantly implicated in motor execution, reward processing, executive control, affective regulation, and sensorimotor integration—domains known to be disrupted in PD. Specifically, the Precentral Gyrus, a key component of the primary motor cortex, exhibits marked hypometabolism and altered functional connectivity in PD (Wang et al., 2024b). The left medial orbitofrontal cortex is related to decision-making performance in Parkinson’s disease (Kobayakawa et al., 2016). Moreover, compared with normal controls, patients with PD showed a globally reduced structural-functional connectivity decoupling in the Inferior and Medial Orbitofrontal Cortices (ORBinf.R and ORBsupmed.L) (Zou et al., 2024). The Amygdala, central to emotional processing and reward evaluation, demonstrates reduced volume and disrupted connectivity within limbic-striatal circuits in PD, which may underlie affective dysregulation and apathy (Yoshimura et al., 2005). The Thalamus, a vital relay hub within cortico-basal ganglia-thalamo-cortical loops, is frequently implicated in PD pathophysiology due to its role in motor and non-motor symptomatology, including tremor and cognitive fluctuations (Zirh et al., 1998). The Triangular and Opercular parts of the Inferior Frontal Gyrus (IFGtriang.R and IFGoperc.R), critical for executive control and language processing, demonstrate impaired activation and connectivity in PD, aligning with cognitive inflexibility and verbal fluency deficits (Li et al., 2020). It is worth noting that although there is currently no direct literature indicating a pathological association between the superior temporal gyrus and PD, its involvement in auditory processing and social cognition may indirectly contribute to communication difficulties and social withdrawal symptoms in patients with Parkinson’s disease. Crucially, unlike ADNI and ASD, the ”core hubs” in the interaction network—PCG.R, STG.R, IFGtriang.R, and ANG.R—did not rank among the top positions in the ”key brain importance” list. Conversely, the seven regions listed as ”key brain importance,” such as ACG.R, TPOmid.R, and CAU.L, were almost entirely absent from the top 20 strongest interactions. Together, these findings highlight two distinct perspectives on ”importance” in brain: ”global centrality” versus ”local strong coupling.”
@@
Autism Spectrum Disorder (ASD). As illustrated in Fig. 14, the ten ROIs with the highest attention scores are: Right Inferior Temporal Gyrus (ITG.R), Left Supramarginal Gyrus (SMG.L), Left Inferior Temporal Gyrus (ITG.L), Left Triangular Part of the Inferior Frontal Gyrus (IFGtriang.L), Left Caudate (CAU.L), Right Caudate (CAU.R), Left Middle Temporal Gyrus (TPOmid.L), Right Middle Temporal Gyrus (TPOmid.R), Right Anterior Cingulate Gyrus (ACG.R), and Right Rolandic Operculum (ROL.R). These regions are predominantly implicated in social cognition, language processing, executive control, and sensorimotor integration. Specifically, the Inferior Temporal Gyrus has been repeatedly reported to exhibit atypical activation and connectivity patterns in individuals with ASD, reflecting its crucial role in face perception and semantic processing (Kim et al., 2021). The Supramarginal Gyrus, a core component of the mirror-neuron system, has also been shown to demonstrate aberrant functional connectivity in ASD, which may contribute to deficits in empathy and social interaction (Bi et al., 2025). The Caudate nucleus, given its involvement in reward processing and cognitive flexibility, has been associated with restricted and repetitive behaviors characteristic of ASD (Adorjan et al., 2017). Furthermore, the Middle Temporal Gyrus and the Anterior Cingulate Gyrus have been implicated in the pathophysiology of ASD through their roles in theory-of-mind and conflict monitoring, respectively (Zhu et al., 2022; Xu et al., 2019). The Rolandic Operculum, although less frequently discussed, is closely linked to auditory and sensorimotor integration, the disruption of which may underlie sensory hypersensitivity commonly observed in ASD (Zhang et al., 2025). Notably, the majority of the top ten ROI interactions with the highest attention scores center on ITG.R, suggesting that this region may also function as a pivotal hub within the brain network and is selectively disrupted during the progression of ASD. This finding aligns with recent literature highlighting the central role of the ITG in the neurodevelopmental trajectory of autism (Kim et al., 2021).
@@ 4.4.4 The higher-order organization that differ between patients and healthy individuals
Next, we further investigate whether the HOIs among key brain regions identified by HOI-Brain show significant differences during disease progression, which could offer additional therapeutic insights for brain disease research. For ease of calculation, we selected four out of the top ten key brain regions identified for each disease. Using our proposed Multiplication of Temporal Derivatives (MTD)-based co-fluctuation calculation method, we computed the quadruplet interactions among the relevant brain regions and subsequently compared the group differences. As shown in Fig. 15, 16, 17, compared to other brain regions, although these brain regions all demonstrate significant importance in the diagnosis of AD or ASD—indicating that they may have sustained varying degrees of damage—their HOIs exhibit notable differences across various stages of disease progression. This suggests a dissociation between HOIs in the brain and the functions of the individual brain regions themselves, which is consistent with the literature (Wang et al., 2019). In addition, we observed that during the transition from healthy individuals to those with AD or ASD, the positive HOIs between relevant brain regions gradually weaken, indicating a progressive impairment of higher-order synergistic functions among brain regions. On the other hand, the negative HOIs between these brain regions gradually strengthen, reflecting an increasing degree of internal functional disruption in the brain. This phenomenon is consistent with findings reported in previous literature (Santoro et al., 2024). However, this phenomenon is exactly the opposite in PD, which may be due to the fact that Parkinson’s patients, because of their motor impairments, are prone to hallucinations. Hallucinations in Parkinson’s disease may result from excessive influence of higher-order brain regions on early sensory processing, which is consistent with the enhanced functional integration between sensory and higher-order networks (Tan et al., 2023).
@@
Alzheimer’s disease (AD). As shown in Fig. 15, in particular, the quadruplet interactions among CAU.R, HIP.L, PHG.R, and AMYG.L; the quadruplet interactions among OLF.L, HIP.L, PHG.R, and AMYG.L; and the quadruplet interactions among CAU.R, OLF.L, PHG.R, and AMYG.L, which involve the cortico-striatal-olfactory circuit, exhibit significant inter-group differences during the progression from cognitively normal (CN) to MCI, with p-values <0.05<0.05< 0.05. However, these differences become less pronounced during the progression from MCI to AD, with p-values 0.05$\geq$ 0.05≥ 0.05. This suggests that these HOIs may serve as potential biomarkers for the early diagnosis of Alzheimer’s disease. In contrast, the quadruplet interactions among CAU.R, TPOmid.L, HIP.L, and AMYG.L, which involves the ’social-reward-emotion-memory’ integrated circuit, shows significant inter-group differences throughout the entire AD progression, with p-values <0.001<0.001< 0.001. This indicates that these HOIs may be persistently disrupted across the entire disease course, playing a critical role in AD pathogenesis.
@@
Parkinson’s disease (PD). As shown in Fig. 16, the quadruplet interactions among PCG.R, ORBsupmed.L, IPL.R, and SPG.R; the interactions among PCG.R, ORBsupmed.L, IPL.R, and IFGoperc.R; and those among PCG.R, ORBsupmed.L, IPL.R, and STG.R, which collectively engage the dorsal attention network and fronto-parietal executive control circuits, exhibit significant inter-group differences ($p<0.05$) between healthy controls and early-stage PD patients. These findings suggest that these functional connectivity patterns may serve as potential early biomarkers for cognitive-motor integration deficits characteristic of the prodromal stages of PD. In addition, the quadruplet interactions involving PCG.R, ORBsupmed.L, THA.R, and IFGoperc.R, which map onto the mesocortical limbic-striatal-thalamic loop, show inter-group differences ($p<0.001$). This persistent disruption aligns with the progressive degeneration of dopaminergic pathways and thalamocortical relay integrity, implicating these interactions in both motor dysfunction (e.g., bradykinesia) and nonmotor symptoms (e.g., apathy, executive dysfunction) throughout the pathogenesis of PD-related motor and cognitive impairment.
@@
Autism Spectrum Disorder (ASD). From Fig. 17, we observed that the quadruplet interactions among the top five key brain regions—including the interactions between ITG.R, SMG.L, ITG.L, and IFGtriang.L; between SMG.L, ITG.L, IFGtriang.L, and CAU.L; between ITG.R, ITG.L, IFGtriang.L, and CAU.L; between ITG.R, SMG.L, IFGtriang.L, and CAU.L; and between ITG.R, SMG.L, ITG.L, and CAU.L, which involves a ”visual-auditory-semantic-action-control” closed-loop system. These interactions exhibited significant inter-group differences throughout the entire progression of ASD. Such quadruplet interactions may serve as potential biomarkers for investigating tool-use deficits, semantic control impairments, and action-language disconnection in social communication associated with autism. However, the quadruplet interactions among ITG.L, IFGtriang.L, CAU.L, and CAU.R show no significant inter-group differences throughout the entire AD progression, with p-values 0.05$\geq$ 0.05≥ 0.05. This suggests a dissociation between higher-order brain interactions and the functions of individual brain regions themselves.
@@ 5.1 The model
In this study, we focused on the importance of signed higher-order (group) interactions in fMRI signals, which were inferred using a recently developed topological approach (Hyekyoung et al., 2011). Our proposed HOI-Brain incorporates: (i) a novel higher-order interaction representation strategy via Multiplication of Temporal Derivatives (MTD) to quantify dynamic functional co-fluctuations of group ROIs; (ii) a novel higher-order feature extraction method that links HOIs in the brain to higher-order organizational structures, focusing on quadruplet-level interaction signatures and two-dimensional void descriptors extracted through Persistent Homology; (iii) a pioneering effort to distinguish between positively and negatively synergistic higher-order interactions in the brain; (iv) a novel multi-channel brain network Transformer that integrates lower-order edge features with higher-order topological invariants. This new framework emphasizes the potential of quadruplet or higher-order region interactions — representing critical advancements for precision medicine, a deeper understanding of neurological disorders, and broader contributions to neuroimaging research. It demonstrates superior prediction accuracy for classifying Alzheimer’s Disease, Parkinson’s Disease, and Autism Spectrum Disorder compared to traditional machine learning models, graph neural network (GNN)-based models, Transformer-based models, hypergraph neural network (HGNN)-based models, and the Persistent Homology (PH)-based models. Additionally, ablation study also indicates that the effectiveness of quadruplet interactions compared to triplet interactions, the effectiveness of distinguishing between positively and negatively synergistic HOIs, and the effectiveness of different components in model.
@@
Furthermore, an investigation was conducted into the most discriminative neural patterns using an attention mechanism. This investigation revealed that quadruplet-level interaction signatures exhibit substantially greater importance than two-dimensional void descriptors and edge features in both the ADNI and ABIDE datasets. It is noteworthy that negative synergistic quadruplets tend to exhibit a higher degree of significance in comparison to their positive synergistic counterparts. Furthermore, positive synergistic voids have been observed to exceed negative synergistic voids in terms of their impact. This novel observation may offer critical insights into the pathological mechanisms underlying brain disorders. In order to provide further validation of these findings, inter-group comparisons were performed of signed higher-order topological signatures in the ADNI dataset. The findings demonstrated a substantial discrepancy in the quantity of quadruplet-level interaction signatures among the groups ($p<0.001$) in comparison to void descriptors. Furthermore, positive voids demonstrated more pronounced inter-group disparities than negative voids. The findings suggest that disease progression may primarily disrupt higher-order neural interactions, which subsequently influence high-dimensional hole structures. An intriguing pattern emerged in the comparative analysis of CN individuals and those with AD. CN individuals appear to rely more on positive information for whole-brain topological organisation, whereas AD patients show a greater dependence on negative information. Furthermore, longitudinal trends indicate that the number of positive quadruplets gradually declines with disease progression, while negative quadruplets follow a biphasic trajectory – decreasing initially and then increasing. This suggests that during the transition from CN to MCI, there is a decline in both positive and negative synergistic interactions. However, as the condition progresses from MCI to AD, there is an increasing adoption of negative synergy by neural networks, alongside a suppression of excitability. This may be a compensatory mechanism employed by the brain to preserve functional integrity across neural modules.
@@
In the context of investigating the significant brain regions and their interactions in the context of brain diseases, the utilisation of HOI-Brain facilitates the identification of these regions from a more comprehensive perspective, a departure from the low-level perspective that has been adopted in previous studies (Yan et al., 2019; Li et al., 2021). The findings, which function as potential biomarkers, are consistent with a substantial corpus of extant literature (Zidan et al., 2019; Zhi et al., 2024; Ye et al., 2019; Kim et al., 2021). Furthermore, they furnish novel insights with regard to the diagnosis of brain diseases. The right caudate has been identified as a potential hub node within the brain network. Previous literature has indicated that disruptions in this region are consistent with the progression of Alzheimer’s disease (Zhi et al., 2024). In a similar manner, the ITG.R has been observed to function as an additional pivotal hub within the brain network. It has been demonstrated that this region is selectively disrupted during the progression of ASD (Kim et al., 2021).